A force (-4i+2j)AND moves an object from the point p(1,3)m to the point Q(3,8) . the work done by the force is


1
Expert's answer
2021-05-30T13:23:16-0400

By the definition of the work done, we have:


"W=Fd."

Let's first find the displacement of the object:


"d=(3-1)i+(8-3)j=(2i+5j)\ m."

Finall, we can find the work done by the force:


"W=(-4i+2j)\ N\cdot(2i+5j)\ m,""W=-4 N⋅2 m+2 N⋅5 m=2\ J."